# Quantum-impurity sensing of altermagnetic order

### v2

@@ Abstract
Introduction – Altermagnetism is a recently identified magnetic phase characterized by a unique spin arrangement of electrons in solids [1, 2, 3, 4]. Like antiferromagnets (AFMs), altermagnets (ALMs) exhibit vanishing net magnetization. However, unlike AFMs, ALMs exhibit time-reversal broken and spin-polarised bands, a characteristic of ferromagnets. Unlike the s-wave-like isotropic spin polarisation in ferromagnets, the spin polarisation in ALMs takes d, g, or i-wave form, i.e. a momentum-space structure of spin-split bands with energy iso-surfaces that are anisotropic, and intersect at 2, 4, or 6 spin-degenerate nodes [1]. Altermagnetic order was proposed in a wide range of materials [2, 5], including metals and insulators, and two dimensional systems [6], and it has been experimental confirmed in a few materials [7, 8, 9, 10] and in some cases remains stable under ambient conditions [9, 11, 12]. Its discovery motivated by predicition and observation of unconventional crystal anomalous Hall effect [13, 14, 15, 16] opens up exciting research avenues [1, 2, 5, 3], including magnetoresistance effects [17], topological [6, 18], multiferroic properties [19, 20, 21], and interplay with superconductivity [22], to name a few.
@@
Figure 1: (a) Relaxometry of spin diffusion in altermagnets. The relaxation rate $\Gamma$ of a quantum impurity with transition frequency $\omega$ is monitored for different relative orientations, described by the set of angles $\{\Phi\}=(\theta,\varphi)$, and at different distances $d$ to the altermagnet. The spin diffusion length $l_{0}$ sets the relevant distance scale. (b) The relaxation contrast over a set of relative orientations exhibits a non-trivial dependence with the distance, which is a fingerprint unique to altermagnetism. Its value can increase up to $\sim 27$ %.
@@
The stray magnetic field correlations carry information about single and multiple magnon scattering processes in the sensed magnet [33, 34, 29, 35, 32, 36, 31]. For a NV with frequency $\omega$ lying within the magnon gap, single magnon processes will be off resonant and thus will not contribute to the relaxation rate. In this case, the main contribution to $\Gamma$ is the longitudinal correlation function $\langle\{\hat{s}_{\parallel}(\vec{r},t),\hat{s}_{\parallel}(\vec{r}^{\prime},0)\}\rangle$, where $\parallel$ refers to the component of the spin density operator $\hat{\vec{s}}$ parallel to the Néel vector. This correlation function is associated with two magnon processes and, thus, with spin transport. These conditions hold for a NV-vacancy center with $\omega\sim$ GHz [37], and an ALM in a weak bias field and with easy-axis anisotropy [38, 39, 40, 41].
@@
$\Gamma[\omega]=\frac{\hbar\gamma^{2}\tilde{\gamma}^{2}}{2}{\rm{coth}}\left(\frac{\hbar\beta\omega}{2}\right)\int\frac{d^{2}\vec{k}}{(2\pi)^{2}}\mathcal{C}_{\{\Phi\}}(d,\vec{k})\chi^{\prime\prime}_{\parallel}(\omega,\vec{k}),$ (2)
@@
where $\tilde{\gamma}$ is the magnet’s gyromagnetic ratio and $d$ is the QI-sample distance. The geometric factor $\mathcal{C}_{\{\Phi\}}(d,\vec{k})=(2\pi)^{2}k^{2}e^{-2kd}\mathcal{F}_{\{\Phi\}}(\phi_{k})$ (with ${\rm{tan}}(\phi_{k})=k_{y}/k_{x}$) is obtained from the magnetostatic Green’s function of a dipole [43], which is commonly encountered in NV-based noise spectroscopy [34, 32, 36], and depends on the relative orientation between the NV principal axis (assumed to be in the $yz$ plane) and the in-plane Néel vector of the ALM via the function $\mathcal{F}_{\{\Phi\}}(\phi_{k})$, where the set of angles $\{\Phi\}=(\theta,\phi)$ describe the orientation of the NV principal axis and the Néel vector, as illustrated in Fig. 1. We remark that our formalism would also be suitable for Néel vectors with out-of-plane components. The coefficient $k^{2}e^{-2kd}$ in $\mathcal{C}_{\{\Phi\}}(d,\vec{k})$ introduces a momentum filter function: for $k\gtrsim 1/d$ the integrand is suppressed by the exponential factor $e^{-2kd}$, and at the same time it dampens it for $k\lesssim 1/d$ due to the term proportional to $k^{2}$. This basically means that we can single out the response function Eq. (2) around $k\simeq 1/d\equiv k_{d}$ in the momentum integral involved in the calculation of the relaxation rate (2), or in physical terms, only magnons with wavelengths $\simeq 1/d$ contribute. The central result of this Letter is that the relaxation rate in Eq. (2) can distinguish between spin diffusion in ALMs and in more conventional magnets such as antiferromagnets (AFMs), as a result of the anisotropy of the diffusion response function $\chi_{\parallel}(\omega,\vec{k})$ in $k$-space – a consequence of the alternating magnon band structure of an ALM [38, 40].
Spin diffusion in ALMs – We obtain the response function $\chi_{\parallel}$ in the diffusive limit following the phenomenological transport approach in [44, 45]. ALMs can be modeled through two magnetic sub-lattices, and thus host two species of magnons, indicated by $\alpha$ and $\beta$, which carry opposite spins [45]. When the system is perturbed away from thermal equilibrium, for example, due to a temperature gradient, or a magnetic field [46], magnons can diffuse, thus transporting spin. For each of the two ALM magnon species, we assume that the excess magnon density with respect to its equilibrium value, which we denoted by $\delta n_{\xi=\alpha,\beta}$, behaves diffusively:
$\partial_{t}\delta n_{\xi}+\nabla\cdot\vec{j}_{\xi}$ $=-\frac{1}{\tau_{s}}(\delta n_{\xi}\pm\chi_{0}h),$ (3)
where $h$ is a generalized force thermodynamically conjugated to $\delta n_{\xi}$, $\tau_{s}$ is the relaxation time, and $\vec{j}_{\xi}=-\overleftrightarrow{D}_{\xi}\cdot\nabla(\delta n_{\xi}\pm\chi_{0}h)$ is the magnon density current, which is given in terms of the diffusion tensor $\overleftrightarrow{D}_{\xi}$. The static spin susceptibility is indicated by $\chi_{0}$, and the associated chemical potential is $(\delta n_{\xi}/\chi_{0}\pm h)$ [47], with the different signs corresponding to the different magnon species [46]. Relying on a Boltzmann distribution for the steady-state magnon occupation [45], we can obtain the elements of the diffusion tensor [48], which depend on the magnon band structure, and the magnon momentum relaxation rate [41].
@@
As a simple way to model magnon bands in a 2D ALM [49, 50, 51], we resort to the Lieb lattice of Fig. 2(a) with the Heisenberg Hamiltonian
@@
The spin diffusion response function is then obtained from the set of coupled linear diffusion equations for $\mathcal{S}_{\parallel}=\hbar(-\delta n_{\alpha}+\delta n_{\beta})$, related to the longitudinal spin density, and $\mathcal{S}_{\perp}=\hbar(\delta n_{\alpha}+\delta n_{\beta})$, related to the total magnon density. The dynamics of the magnon densities are given by Eq. (3), and the diffusion tensor couples $\mathcal{S}_{\parallel}$ to $\mathcal{S}_{\perp}$. These equations can be solved in Fourier domain through $\mathcal{S}_{\parallel}(\omega,\vec{k})=\chi_{\parallel}(\omega,\vec{k})h$, with the response function is given by [48]
$\chi_{\parallel}(\omega,\vec{k})=\frac{\hbar\chi_{0}\left[D_{1}k^{2}+\frac{1}{\tau_{s}}-\frac{D_{2}^{2}k^{4}\cos^{2}(2\phi_{k})}{-i\omega+D_{1}k^{2}+1/\tau_{s}}\right]}{-i\omega+D_{1}k^{2}+\frac{1}{\tau_{s}}-\frac{D_{2}^{2}k^{4}\cos^{2}(2\phi_{k})}{-i\omega+D_{1}k^{2}+1/\tau_{s}}}.$ (5)
The dependence of $\chi_{\parallel}$ on the directionality in $k$-space and the factor $\propto k^{4}$ are traits unique of ALMs, not present in usual magnets [29, 30, 31].
Relaxation rate and contrast – The response function $\chi_{\parallel}$ in Eq. (5) is anisotropic in $k$-space and includes the dependence $k^{4}$. As a consequence, the dependence of the relaxation rate on the NV-sample distance $d$ and on the relative orientation $\{\Phi\}$ for ALMs is different than that of conventional magnets. To better highlight this feature, we find it useful to define the relaxation contrast
@@
$\Gamma[\omega]|_{d\gg l_{0}}\propto\frac{l_{0}^{4}}{d^{4}}\int_{0}^{2\pi}d\phi_{k}\mathcal{F}_{\theta;\vartheta,\varphi}(\phi_{k}),$ (7)
where we have defined $l_{0}=\sqrt{D_{1}\tau_{s}}$, the isotropic spin diffusion length. Such a relaxation rate has exactly the same form one would get from a magnet with isotropic magnon bands [29, 30]. The value of the contrast is independent on whether the material is an ALM or an AFM, a consequence of the anisotropy of the magnetic noise driving the QI with respect to the orientation angle $\varphi$.. The Néel vector sets the direction of the longitudinal modes probed by relaxometry, rendering the relaxation rate intrinsically anisotropic. ALMs induces further anisotropy contributions due to the magnon splitted bands, which, according to Eq. (5), is washed away for wave-vectors $k\ll 1/l_{0}$. Thus, the NV is insensitive to such additonal trait intrinsic to ALMs when placed at distances $d\gg l_{0}$.
@@
$\Gamma[\omega]|_{d\ll l_{0}}\propto\frac{D_{1}^{2}l_{0}^{2}}{d^{2}}\int_{0}^{2\pi}d\phi_{k}\frac{\mathcal{F}_{\theta;\vartheta,\varphi}(\phi_{k})}{D_{1}^{2}-D_{2}^{2}\cos^{2}(2\phi_{k})}.$ (8)
The anisotropy of the response function Eq. (5) induces an additional dependence of the integrand from directionality in $k$-space in the above equation. In turn, the contrast in this limit is given by
@@
In the case of an AFM, i.e. by setting $D_{2}=0$, the contrast yields $11/21$, which is the same value it has at distances $d\gg l_{0}$. For an ALM, the contrast behaves non-trivially with distance, exhibiting a larger value at closer distances. For $D_{2}\approx D_{1}$ (its maximum value), the contrast at closer distances is $8/12\approx 0.67$, which is an increase of 2727% compared to the large distance limit. This is the case shown in Fig. 1(b), for which we took $D_{2}\approx 1$. In general, the contrast depends on the ratios $D_{2}/D_{1}$ and $d/l_{0}$, and on the product $\omega\tau_{s}$ [48], thus its behavior with respect to the other parameters can also be exploited for extracting intrinsic diffusion parameters. We would like to emphasize that the behavior of the contrast showed here is a consequence of the anisotropy of the response function in Eq.  (5), generated by the off-diagonal component of the diffusion tensor $D_{2}$.
We show in Fig. 3 the relaxation rate $\Gamma$ and the contrast $\mathcal{C}$ in as a function of the distance $d$ for different values of the parameter $\Delta$. The relaxation rate is shown in 3(a) in units of the characteristic rate $\Gamma_{c}=\hbar\chi_{0}\gamma^{2}\tilde{\gamma}^{2}k_{b}T\tau_{s}/l_{0}^{4}$. The relaxation rate for an ALM, depicted by the dark curves, exhibits a different curvature at the cross-over region $d=l_{0}$ when compared to an usual AFM, depicted in light blue, due to the term $\propto k^{4}$ appearing in the susceptibility Eq. (5). Furthermore, as depicted in Fig. 3(b), the contrast increases as the distance $d$ is decreased, a trait that is unique to ALMs and is a consequence of the anisotropy of the spin diffusion susceptibility.
@@
An alternative approach to probing ALMs is through $T_{2}$ relaxometry [63]. The typically stronger $T_{2}$ signal, as compared to $T_{1}$ noise, together with its operational low-frequency range, makes it well-suited for studying systems with low-energy and gapless modes. These include, the longitudinal diffusive modes, hydrodynamic sound modes, as well as nodal magnons in MALMs mentioned above. Furthermore, recent advances in NV-based sensing allow for multi-qubit noise spectroscopy, which can provide spatially resolved information on magnetic noise correlations [64, 28]. This technique offers powerful means to probe non-invasively spatio-temporal spreading of correlations in ALMs [65].
Acknowledgments - H.H. and J.M. are grateful to E. Demler, P. Dolgirev, S. Gopalakrishnan for related discussions on NV sensing. VASVB acknowledges financial support from the Contrat Triennal 2021-2023 Strasbourg Capitale Europeenne. JS and LS acknowledge support by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - DFG (Project 452301518) and TRR 288 – 422213477 (project A09 and A12). LS acknowledges support by the ERC Starting Grant No. 101165122. JM acknowledges support by the DFG through the grant HADEQUAM-MA7003/3-1. We acknowledge support from the Dynamics and Topology Centre, funded by the State of Rhineland Palatinate.

### v3

@@ Abstract
Introduction – Altermagnetism is a recently identified magnetic phase characterized by a unique spin arrangement of electrons in solids [1, 2, 3, 4]. Like antiferromagnets (AFMs), altermagnets (ALMs) exhibit vanishing net magnetization. However, unlike AFMs, ALMs exhibit time-reversal broken and spin-polarised bands, a characteristic of ferromagnets. Unlike the s-wave-like isotropic spin polarisation in ferromagnets, the spin polarisation in ALMs takes d, g, or i-wave form, i.e. a momentum-space structure of spin-split bands with energy iso-surfaces that are anisotropic, and intersect at 2, 4, or 6 spin-degenerate nodes [1]. Altermagnetic order was proposed in a wide range of materials [2, 5], including metals and insulators, and two dimensional systems [6], and it has been experimental confirmed in a few materials [7, 8, 9, 10] and in some cases remains stable under ambient conditions [9, 11, 12]. Its discovery motivated by predicition and observation of unconventional crystal anomalous Hall effect [13, 14, 15, 16] opens up exciting research avenues [1, 2, 5, 3], including magnetoresistance effects [17], topological [6, 18, 19], multiferroic and ferroelectric properties [20, 21, 22, 23, 24, 25], spin and exciton dynamics [26, 27], and interplay with superconductivity [28], to name a few.
@@
Figure 1: (a) Relaxometry of spin diffusion in altermagnets. The relaxation rate $\Gamma$ of a quantum impurity with transition frequency $\omega$ is monitored for different relative orientations, described by the angles $(\theta,\varphi)$, and at different distances $d$ to the altermagnet. The spin diffusion length $l_{0}$ sets the relevant distance scale. (b) The relaxation contrast over a set of relative orientations exhibits a non-trivial dependence with the distance, which is a fingerprint unique to altermagnetism. Its value can increase up to $\sim 27$ %.
@@
The stray magnetic field correlations carry information about single and multiple magnon scattering processes in the sensed magnet [40, 41, 38, 42, 39, 43, 37]. The relaxation rate, thus depends on both transversal and longitudinal spin correlations [38, 37]. The transversal correlations are associated with single magnon processes, which are relevant if the frequency, $\omega$, of the NV is at resonance with the magnon frequency (plus a broadening due to magnon relaxation). For altermagnets, the typical magnon gap is $\sim$ THz [44], while the NV frequency is $\sim$ GHz, and thus contributions due to transverse spin correlations to the NV relaxation are negligible. In this case, the main contribution to $\Gamma$ is the longitudinal correlation function $\langle\{\hat{s}_{\parallel}(\vec{r},t),\hat{s}_{\parallel}(\vec{r}^{\prime},0)\}\rangle$, where $\parallel$ refers to the component of the spin density operator $\hat{\vec{s}}$ parallel to the Néel vector. This correlation function is associated with two magnon processes and, thus, with spin transport. These conditions hold for a NV center with $\omega\sim$ GHz [45], and an ALM in a weak bias field and with easy-axis anisotropy [46, 47, 44, 48].
@@
$\Gamma[\omega]=\frac{\hbar\gamma^{2}\tilde{\gamma}^{2}}{2}{\rm{coth}}\left(\frac{\hbar\beta\omega}{2}\right)\int\frac{d^{2}\vec{k}}{(2\pi)^{2}}\mathcal{C}_{\theta,{\color[rgb]{0,0,0}\varphi}}(d,\vec{k})\chi^{\prime\prime}_{\parallel}(\omega,\vec{k}),$ (2)
@@
where $\tilde{\gamma}$ is the magnet’s gyromagnetic ratio,$d$ is the QI-sample distance, and $\beta=1/k_{B}T$, where $T$ is the temperature and $k_{B}$ is the Boltzmann constant. The geometric factor $\mathcal{C}_{\theta,{\color[rgb]{0,0,0}\varphi}}(d,\vec{k})=(2\pi)^{2}k^{2}e^{-2kd}\mathcal{F}_{\theta,{\color[rgb]{0,0,0}\varphi}}(\phi_{k})$ (with ${\rm{tan}}(\phi_{k})=k_{x}/k_{z}$) is obtained from the magnetostatic Green’s function of a dipole [50], which is commonly encountered in NV-based noise spectroscopy [41, 39, 43], and depends on the relative orientation between the NV principal axis (assumed to be in the yzyz plane) and the in-plane Néel vector of the ALM via the function $\mathcal{F}_{\theta,{\color[rgb]{0,0,0}\varphi}}(\phi_{k})$, where the angles $(\theta,{\color[rgb]{0,0,0}\varphi})$ describe the orientation of the NV principal axis and the Néel vector, as illustrated in Fig. 1. We remark that our formalism would also be suitable for Néel vectors with out-of-plane components. The coefficient $k^{2}e^{-2kd}$ in $\mathcal{C}_{\theta,{\color[rgb]{0,0,0}\varphi}}(d,\vec{k})$ , which is a consequence of the dipolar nature of the interaction between the NV spin and the magnetic excitations, introduces a momentum filter function: for $k\gtrsim 1/d$ the integrand is suppressed by the exponential factor $e^{-2kd}$, and at the same time it dampens it for $k\lesssim 1/d$ due to the term proportional to $k^{2}$. This means that we can single out the response function Eq. (2) around $k\simeq 1/d\equiv k_{d}$ in the momentum integral involved in the calculation of the relaxation rate (2), or in physical terms, only magnons with wavelengths $\simeq 1/d$ contribute. The central result of this Letter is that the relaxation rate in Eq. (2) can distinguish between spin diffusion in ALMs and in more conventional magnets such as antiferromagnets (AFMs), as a result of the anisotropy of the diffusion response function $\chi_{\parallel}(\omega,\vec{k})$ in $k$-space – a consequence of the alternating magnon band structure of an ALM [46, 44].
Spin diffusion in ALMs – We obtain the response function $\chi_{\parallel}$ in the diffusive limit following the phenomenological transport approach in [51, 52] adapted to ALMs [53]. We assume that the magnon excess density $\delta n_{\alpha,\beta}$ (for $\alpha$ and $\beta$ the two magnon species) follows a diffusion equation, with a diffusion tensor $\overleftrightarrow{D}_{\alpha,\beta}$ that depends on the magnon band structure, and the magnon momentum relaxation rate [48, 53].
@@
As a simple way to model magnon bands in a 2D ALM [54, 55, 56, 57, 58], we resort to the Lieb lattice of Fig. 2(a) with the Heisenberg Hamiltonian
@@
The spin diffusion response function is then obtained from the set of coupled linear diffusion equations for $\mathcal{S}_{\parallel}=\hbar(-\delta n_{\alpha}+\delta n_{\beta})$, related to the longitudinal spin density, and $\mathcal{S}_{\perp}=\hbar(\delta n_{\alpha}+\delta n_{\beta})$, related to the total magnon density [53]. These equations can be solved in Fourier domain through $\mathcal{S}_{\parallel}(\omega,\vec{k})=\chi_{\parallel}(\omega,\vec{k})h$, with the response function given by [53]
$\chi_{\parallel}(\omega,\vec{k})=\frac{\hbar\chi_{0}\left[D_{1}k^{2}+\frac{1}{\tau_{s}}-\frac{D_{2}^{2}k^{4}\cos^{2}(2\phi_{k})}{-i\omega+D_{1}k^{2}+1/\tau_{s}}\right]}{-i\omega+D_{1}k^{2}+\frac{1}{\tau_{s}}-\frac{D_{2}^{2}k^{4}\cos^{2}(2\phi_{k})}{-i\omega+D_{1}k^{2}+1/\tau_{s}}},$ (4)
where $\tau_{s}$ is the spin relaxation time. The dependence of $\chi_{\parallel}$ on the directionality in $k$-space due to factor $\propto k^{4}\cos^{2}(2\phi_{k})$ is a unique trait of ALMs, not present in usual magnets [38, 36, 37].
Relaxation rate and contrast – The response function $\chi_{\parallel}$ in Eq. (4) is anisotropic in $k$-space. As a consequence, the dependence of the relaxation rate on the relative orientation $(\theta,\varphi)$ at a given distance $d$ is different for ALMs when compared with conventional magnets. To better highlight this feature, we find it useful to define the relaxation contrast
@@
$\Gamma[\omega]|_{d\gg l_{0}}\propto\frac{l_{0}^{4}}{d^{4}}\int_{0}^{2\pi}d\phi_{k}\mathcal{F}_{\theta,\varphi}(\phi_{k}),$ (6)
where we have defined $l_{0}=\sqrt{D_{1}\tau_{s}}$, the isotropic spin diffusion length. Such a relaxation rate has exactly the same form one would get from a magnet with isotropic magnon bands [38, 36]. The value of the contrast is independent on whether the material is an ALM or an AFM, a consequence of the anisotropy of the magnetic noise driving the QI with respect to the orientation angle $\varphi$. The Néel vector sets the direction of the longitudinal modes probed by relaxometry, rendering the relaxation rate intrinsically anisotropic. ALMs induces further anisotropy contributions due to the magnon splitted bands, which, according to Eq. (4), is washed away for wave-vectors $k\ll 1/l_{0}$. Thus, the NV is insensitive to such additonal trait intrinsic to ALMs when placed at distances $d\gg l_{0}$.
@@
$\Gamma[\omega]|_{d\ll l_{0}}\propto\frac{D_{1}^{2}l_{0}^{2}}{d^{2}}\int_{0}^{2\pi}d\phi_{k}\frac{\mathcal{F}_{\theta,\varphi}(\phi_{k})}{D_{1}^{2}-D_{2}^{2}\cos^{2}(2\phi_{k})}.$ (7)
While the relaxation rate exhibits the same scaling $d^{-2}$ as AFMs, the anisotropy of the response function Eq. (4) induces an additional dependence of the integrand from directionality in $k$-space in the above equation. In turn, the contrast in this limit is given by
@@
In the case of an AFM, i.e. in the limit $D_{2}\rightarrow 0$ with $D_{1}$ finite, the contrast yields $11/21$, which is the same value it has at distances $d\gg l_{0}$. For an ALM, the contrast behaves non-trivially with distance, exhibiting a larger value at closer distances. For $D_{2}\approx D_{1}$ (its maximum value), the contrast at closer distances is $8/12\approx 0.67$, which is an increase of 2727% compared to the large distance limit, as shown in Fig. 1(b). In general, the contrast depends on the ratios $D_{2}/D_{1}$ and $d/l_{0}$, and on the product $\omega\tau_{s}$ [53], thus its behavior with respect to the other parameters can also be exploited for extracting intrinsic diffusion parameters. Such a behavior of the contrast is a consequence of the anisotropy of the response function in Eq. (4), generated by the off-diagonal component of the diffusion tensor $D_{2}$. We show in Fig. 3 the relaxation rate $\Gamma$, in units of the characteristic rate $\Gamma_{c}=\hbar\chi_{0}\gamma^{2}\tilde{\gamma}^{2}k_{b}T\tau_{s}/l_{0}^{4}$, and the contrast $\mathcal{C}$ in as a function of the distance $d$ for different values of the parameter $\Delta$. The relaxation rates for ALMs exhibit a behavior with the distance $d$ similar to AFMs, as show in Fig. 3(a). Nevertheless, the contrast, depicted in Fig. 3(b), increases as the distance $d$ is decreased. This is a trait that is unique to ALMs, and a consequence of the anisotropic term of Eq. (4).
@@
The relaxation rate induced by the spin transport should dominate the intrinsic relaxation rate $\Gamma^{(0)}$ experienced by the QI in the absence of magnetic noise. This is particularly relevant for the measurement of the spin diffusion noise at farther distances, for which the spin-diffusion induced relaxation is weaker. $\Gamma^{(0)}$ includes contributions due to coupling to phonons [63, 64], magnetic noise [63, 64], and due to electric and strain noise [65, 66]. The first two sources can be minimized by operating the system at low temperatures and using high-quality diamonds [67]. The electric and strain contributions depend mainly on the distance between the NV center and the surface of the diamond [65, 66], and can be optimized with proper fabrication techniques.
@@
An alternative approach to probing ALMs is through $T_{2}$ relaxometry [74]. The typically stronger $T_{2}$ signal, as compared to $T_{1}$ noise, together with its operational low-frequency range, makes it well-suited for studying systems with low-energy and gapless modes. These include, the longitudinal diffusive modes, hydrodynamic sound modes, as well as nodal magnons in MALMs mentioned above. Recent advances in NV-based sensing allow for multi-qubit noise spectroscopy, which can provide spatially resolved information on magnetic noise correlations [75, 34]. This technique offers powerful means to probe non-invasively spatio-temporal spreading of correlations in ALMs [76].
Acknowledgments - $H.$ and J.M. are grateful to E. Demler, P. Dolgirev, S. Gopalakrishnan for related discussions on NV sensing. VASVB acknowledges financial support from the Contrat Triennal 2021-2023 Strasbourg Capitale Europeenne. JS and LS acknowledge support by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - DFG (Project 452301518) and TRR 288 – 422213477 (project A09 and A12). LS acknowledges support by the ERC Starting Grant No. 101165122. JM acknowledges support by the DFG through the grant HADEQUAM-MA7003/3-1. We acknowledge support from the Dynamics and Topology Centre, funded by the State of Rhineland Palatinate.